OpenStudy (anonymous):

from a point 13 ft from the base of a building the angle of elevation from the ground to the top of the building is 70 degrees which is closed to the height of the bulding ? A) 12ft B)23ft C)36ft D)41ft

OpenStudy (anonymous):

#2

OpenStudy (anonymous):

@LizzyLove<3 can u help me on this one?

OpenStudy (anonymous):

yeah ill try... hold on

OpenStudy (anonymous):

okay thank you :) sorry i just dont get it

OpenStudy (anonymous):

Do u see the right angle and the 70degrees angle??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Add those together, and if u know that a triangle is 180, the missing angle would be 30, right?

OpenStudy (anonymous):

180 degrees*

OpenStudy (anonymous):

oh yes so i add 70 + 90 ?

OpenStudy (anonymous):

yea and u get...?

OpenStudy (anonymous):

160

OpenStudy (anonymous):

160 right?

OpenStudy (anonymous):

okk yes.. good, now take 180 minus 160 and u get what?

OpenStudy (anonymous):

yes so it 20?

OpenStudy (anonymous):

yeas but ur question says which is closest, so the answer would be 23ft u get it?

OpenStudy (anonymous):

oh so i estamiate lol i dont now how to spell it okay wat about that 13 ft we dont use it?

OpenStudy (anonymous):

it's just there to throw you off, like alot of problems are:) and yeas u do estimate, do u need anymore help?

OpenStudy (anonymous):

oh okay :) so i just leave it and yes plz on another question wait

OpenStudy (anonymous):

the angle from a point on the ground to the top of a 168 ft tower is 35 degrees

OpenStudy (anonymous):

ok brb

OpenStudy (anonymous):

a) sin 35 degrees =168/a b)tan 55 degrees= 168/a c) cos 35 =168/a d) cos 55 =a/168

OpenStudy (anonymous):

oh, now i feel bad because idk how to do sin, cos, or tan. i don't even know what they are!!:( SORRY

OpenStudy (anonymous):

oh its okay so can u help me on another ?

OpenStudy (anonymous):

yeah i can try:) Thxs btw

OpenStudy (anonymous):

okay can u see wat i attach wat about #5? if u ant c it i can type it

OpenStudy (anonymous):

ur welcome :D

OpenStudy (anonymous):

yeah i can see wait one min:)

OpenStudy (anonymous):

its C:)

OpenStudy (anonymous):

can u plz explain im sorry i just want 2 know how 2 do it